OpenStudy (anonymous):

help! lim x cos x-x^2 / 2x approaches x-> 0

OpenStudy (misty1212):

\[\frac{x\cos(x)-x^2}{2x}\]?

OpenStudy (anonymous):

yeah is that one mistry :D

OpenStudy (anonymous):

hahah Thanks girls!

OpenStudy (misty1212):

you can factor out an \(x\) and cancel first

OpenStudy (misty1212):

\[\frac{\cos(x)-x}{2}\]

OpenStudy (godlovesme):

@emilyboob misty is just helping lol

OpenStudy (godlovesme):

Bye!

OpenStudy (misty1212):

replace x by 0 and that is done